OpenStudy (wwwhhhaaattt?):

Will Fan and medal... DUE IN A COUPLE MINS 4^-3 x 4^6

OpenStudy (anonymous):

\[-8+4=-4\]so the answer is \[3^{-4}=\frac{1}{3^4}=\frac{1}{81}\]

OpenStudy (anonymous):

you watching gracepoint?

OpenStudy (wwwhhhaaattt?):

No the FSU game but can you answer it

OpenStudy (wwwhhhaaattt?):

@satellite73 PLEASE HELP QUICK

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

dunno if that's right..

OpenStudy (anonymous):

think so, though. :)